OpenStudy (anonymous):

Express the complex number in trigonometric form. -3i

jimthompson5910 (jim_thompson5910):

Think of `-3i` as `0-3i = 0 + (-3i)` the last expression is in the form `a+bi` where `a = 0` and `b = -3`

jimthompson5910 (jim_thompson5910):

You need to find r and theta use \[\Large r = \sqrt{a^2 + b^2}\] \[\Large \theta = \arctan\left(\frac{b}{a}\right)\]

OpenStudy (anonymous):

r = √0^2 + -3^2 = √0 + 9 = √9 = 3

jimthompson5910 (jim_thompson5910):

yep `r = 3`

jimthompson5910 (jim_thompson5910):

basically that says: "the point (0,-3) is 3 units away from the origin" `0+(-3i)` can be plotted on the complex plane as the point (0,-3)

jimthompson5910 (jim_thompson5910):

what would theta be?

OpenStudy (anonymous):

im getting an error for theta..

jimthompson5910 (jim_thompson5910):

yes you'll find that b/a = -3/0 = undefined where is tangent undefined?

OpenStudy (anonymous):

cos(x) = 0 ?

jimthompson5910 (jim_thompson5910):

I guess a better way to find theta is by plotting where 0-3i is located 

jimthompson5910 (jim_thompson5910):

what angle is this?